Baskets (an all form) is reinvesting 65% of it's earnings in projects that provide a ROE of 9%. The expected return on similar risky projects is 14% on the stock market. Given the present policy on the firm it's year-end dividend is now $3 per share. At what price will the stock sell?
1
Expert's answer
2021-09-12T19:29:41-0400

We need to use constant growth model to calculate stock price.

Stock price "(P0)=\frac{D1}{(rs-g)}"

Where

D1 =dividend for next years

rs =expected return 

g=Growth rate 

Growth rate (g) =ROE"\times" Retention ratio

              "=9\%×0.65\\

 =5.85\%"

Dividend for next year"(D1) =\$3(1+0.0585)"

           "=\$3.1755"

Expected return(rs) =14%

  Stock price "(P0) =\frac{3.1755}{(0.14-0.0585)}"

               "=\frac{3.1755}{0.0815}\\

 =38.96"

hence stock price is $38.96